find the equation of the circle for which the line segment joining the points(8,0) and (0,-3) is a diameter
 

Grade:11

2 Answers

Meet
6 years ago
It means these points are the diametric ends of the circle so the equation of the circle is (x-8)(x-0) + (y-0)(y+3)=0 and it is x^2+y^2-8x+3y=0
Arun
6 years ago
If we are given with diameter coordinates (x1, y1) & (x2, y2) then 
equation of circle :
(x –x1) (x –x2) + (y – y1)(y – y2) = 0
similarly
(x –8)(x – 0) + (y – 0)(y + 3) = 0
now you can simplify
 
Regards
Arun (askIITians forum expert)